A mass vibrates through amplitude of 2.0cm in simple harmonic motion with period 1.0s. what distance is moved from the distance of oscillation in 0.4s
1
Expert's answer
2021-08-16T08:38:57-0400

A - amplitude (cm)

T - period (s)

t - time variable (s)

x(t) - displacement (cm)


Comply the equation:


"x(t)=A\sin\bigg(\frac{2\pi}{T}t\bigg),\\\space\\
x(0.4)=2\sin\bigg(\frac{2\pi}{1}·0.4\bigg)=0.088\text{ cm}."